OpenStudy (anonymous):

Explain, in complete sentences, which method you would use to solve the following system of equations and why you chose that method. Show your work and provide the solution to the system. 5x + 3y = –7 4x – 2y = –10

OpenStudy (lilg132):

elimination

OpenStudy (anonymous):

I would use elimination method (because I always use this). Multiplying equation 1 by 4 and equation 2 by 5, then subtracting eq 2 from eq 1. thus getting22y= 22, y=1. then substituting y=1 in eq 1

OpenStudy (lilg132):

multiply top equation by 2 multiply bottom by 3 cancel y variable out to solve x

OpenStudy (lilg132):

5x + 3y = –7 (multiply by 2) 4x – 2y = –10 (multiply by 3) 10x + 6y = -14 12x - 6y = -30 add together cancelling out the y variable giving you 22x = -44 22x / 22 = -44 / 22 x = -2

OpenStudy (lilg132):

now use this to work out y

OpenStudy (lilg132):

can you do the rest?

OpenStudy (anonymous):

yes thank you so much<3